What is the pH of a 1.00 molar solution of NaNO2(aq)? The Ka for HNO2 is 7.0*10^-4

(Please include how you solved the problem. Thanks!)
Answer #1

NaNO2 -----> Na+ + NO2-

NO2- ion undergoes hydrolysis.

NO2- + H2O <-----> HNO2 + OH-

1.0                           0          0       initial

-x                            +x        +x     change

1.0-x                         x          x      equilibrium

Ka*Kb = 1.0*10^-14

Kb = (1.0*10^-14)/(7.0*10^-4) ==> 1.43*10^-11

Kb = [NHO2][OH-]/[NO2-]

1.43*10^-11 = x^2/(1)

x = 3.78*10^-6

pOH = -log(3.78*10^-6)

        = 5.42

pH = 14 - 5.42

     = 8.58